How to obtain relationship for converting between Fahrenheit and Celsius scales from the given graph?

Grade:11

2 Answers

Arun
5 years ago
Dear student
 
There is no image attached here. Please check and repost the question with image.
I will be happy to help you.
chetan jangir
4 years ago
You have not attached any graph with it . but i think if u are asking formula to convert it is simple F-32/180=C/100. More simplified F=C9/5+32. We can Also Derive If you Want.